08 Patients Treated With This​ Drug, 159 Developed The Adverse Reaction Of Nausea. Use A 0.10 Significance Level To Test The Claim That 3​% Of Users Develop Nausea. Does Nausea Appear To Be A Problematic Adverse​ Reaction? Identify The ...
This problem has been solved!

See the answer
Consider a drug that is used to help prevent blood clots in certain patients. In clinical​ trials, among 6108 patients treated with this​ drug, 159 developed the adverse reaction of nausea. Use a 0.10 significance level to test the claim that 3​% of users develop nausea. Does nausea appear to be a problematic adverse​ reaction?
Identify the test statistic for this hypothesis test is?
Identify the P value
Answer #1

Answer:

Given,

Null hypothesis Ho : p = 0.03

Alternative hypothesis Ha : p != 0.03

sample n = 6108

x = 159

p^ = x/n

= 159/6108

= 0.0260

consider,

test statistic z = (p^ - p)/(sqrt(p(1-p)/n))

substitute values

= (0.026 - 0.03)/sqrt(0.03(1-0.03)/6108)

z = -3.11

Corresponding p value is 0.0018709 [since from z table]

p value = 0.0019

Here we observe that p value < alpha(0.1), so we reject Ho, there is sufficient evidence.